# Dexterous Contact-Rich Manipulation via the Contact Trust Region

### v3

@@ 1 Introduction
The apparent discrepancy between these two approaches raises the question of how they are related. In fact, analytically deriving the gradients from differentiable simulators shows that the Taylor expansions they produce and the complementarity constraints ubiquitous in contact-implicit trajectory optimization are fundamentally connected (Fiacco, 1983)(Pang et al., 2023, Example 5): the Taylor expansions are obtained by applying the Implicit Function Theorem to the force-balance and complementarity conditions, which form a subset of the contact dynamics constraints typically included in CITO formulations.
@@
To alleviate these limitations, smoothing-based methods relax contact dynamics by reasoning about the average result of all the contact modes around a nominal state given some distribution of state and action disturbances. Recent works (Suh et al., 2022b, a) have attributed the success and scalability of Reinforcement Learning (RL) to this mechanism. Surprisingly, however, Pang et al. (2023); Zhang et al. (2023); Howell et al. (2023a) show that for contact dynamics simulated by convex optimization, this process can be done without Monte-Carlo by using barrier smoothing which performs interior-point relaxation of the original contact dynamics and produces force from a distance. However, despite occasional successes, naively applying classical control methods such as LQR on smoothed linearization of contact dynamics often leads to catastrophic failures (Shirai et al., 2024a).
@@
We show the efficacy of the proposed CTR and MPC method on two representative contact-rich manipulation problems in Figure 1: i) whole-body manipulation on the bimanual iiwa and cylinder system, and ii) in-hand reorientation on the Allegro hand and cube system. For both systems, we aim to answer the following two questions: i) is the MPC method successful in planning for goals under quasidynamic equations of motion (Section 5), and ii) can the MPC method successfully stabilize quasidynamic plans under true second-order dynamics (Section 6)? Through $2000$ trajectory runs in simulation and $100$ runs on hardware, our results delve into the efficacy and the limitations of our approach, as well as the benefits of CTR over the ellipsoidal one in the context of planning and control.
@@ 2.5 Smoothing of Contact Dynamics
[…]
@@
We note that this force i) has a direction opposing the movement in the tangential plane, ii) has a magnitude that satisfies (9), and iii) lies in the friction cone. In addition, plugging (12) into (11) yields
@@ 3.1 Linear Model over Smoothed Primal and Dual Variables
However, many previous works (Suh et al., 2022a; Howell et al., 2023a; Pang et al., 2023; Shirai et al., 2024b; Posa et al., 2014) have noted that due to the non-smooth nature of contact, it is beneficial to smooth the dynamics where a first-order Taylor approximation would be valid beyond the contact mode to which $(\bar{q},\bar{u})$ belongs (Pang et al., 2023). A first-order Taylor expansion under such a smoothed dynamics model can be written as
@@
If we expand the Taylor-approximated equations and drop the terms above first-order, we are left with i) the original equality conditions over nominal coordinates which must hold due to the nominal values being obtained at optimality and ii) its first-order expansion, which must hold due to the definition of the sensitivity gradients (15). ∎
@@ 4.1 Trajectory Optimization with R-CTR
1 Input: Initial state $q_{0}$, goal state $q_{\text{goal}}$, input trajectory guess $\bar{u}_{0:T-1}$, iterations limit $n_{\mathrm{max}}$;
@@ 4.3 Examples
Revisiting the 1D sphere-box system in Figure 3, we would like to push the box to a goal position $q^{\mathrm{o}}_{\text{goal}}$ by running Algorithm 1. We begin with the nominal configuration $\bar{q}_{0}=(\bar{q}_{0}^{\mathrm{o}},\bar{q}_{0}^{\mathrm{a}})=(-0.02,0.2)$, where the ball nearly touches the box, and a nominal action $^{0}\bar{u}=\bar{q}^{\mathrm{a}}_{0}$. The nominal action333 We use the left superscript to denote how the nominal values change through (i) applying the initial guess heuristics, and (ii) iterations in Algorithm 1. The left superscript is intended to be different from the right subscript which indexes time steps. is depicted by the yellow circle in Figure 6a. Even with the help of dynamics smoothing, the slope at this point is still small and thus provides limited guidance for planning.
@@ 4.4 Model Predictive Control
1 Input: Initial state $q_{0}$, goal state $q_{\text{goal}}$, planning horizon $T$, iterations limit $n_{\mathrm{max}}$, MPC rollout horizon $H$;
@@ 6 Stabilization under Second-Order Dynamics
In this section, we would like to understand how the gap between the CQDC and real physics affect MPC performance in more realistic settings. Specifically, we focus on answering the question: can the MPC controller using the CQDC dynamics perform closed-loop stabilization i) on high-fidelity simulated second-order dynamics, and ii) on hardware?
@@ 6.1 Implementation Heuristics for Second-order Dynamics
Firstly, instead of running MPC as quickly as possible, we compute a longer trajectory (usually 1-2 seconds) using Algorithm 2 for multiple steps ($H>1$) (Algorithm 3), roll out $L_{u}$ in its entirety on the second-order dynamics (Algorithm 3), and repeat. Although we can re-plan at a higher rate by simply replacing the CQDC dynamics ($f$ in Algorithm 2 of Algorithm 2) with $f_{\text{2nd}}$, we observed that doing high-rate MPC without waiting for the system to stabilize is inconsistent with the quasi-dynamic nature of the CQDC model. In practice, we found CQDC-based MPC works better when we let the system stabilize to an equilibrium before re-planning in order to ensure $f$ remains similar to $f_{\text{2nd}}$. See Section A.6 for more details about the effect of $H$ on closed-loop tracking performance.
@@
1 Input: Initial state $q_{0}$, goal state $q_{\text{goal}}$, planning horizon $T$, iterations limit $n_{\mathrm{max}}$, MPC rollout horizon $H$, number of re-plans NN;
@@ 7 Generating Goal-conditioned Contact Configurations
To efficiently conduct long-horizon planning for contact-rich manipulation, we utilize two key insights. The first insight comes from dynamic programming: knowledge of key intermediate states allows us to decompose a difficult long-horizon problem into multiple shorter subproblems in which greedy strategies can be more effective. For the planar-pushing example in Figure 16, consider the configuration in which the ball is at the bottom of the box as opposed to the top. If the planner had i) seen this configuration before and ii) known that this configuration is significantly more advantageous for pushing the box upward using local control, it can attempt to go towards this advantageous configuration first before attempting to establish contact with the box.
@@ 7.1 Problem Specification
[…]
@@ 7.1.2 Robustness Regularizer
To evaluate this quantity, we first make a polytopic approximation of the wrench set $\mathcal{W}_{\mathbf{\Sigma},\kappa}^{\mathcal{A}}$ by sampling from the RA-CTR using a rejection sampling scheme. This scheme i) first samples from the ETR ellipsoid and ii) rejects samples that do not obey feasibility constraints (this is the same procedure which generated the samples in Figure 5). Then, we fit a convex hull to these samples using the Qhull library.
@@ 7.2 Solving (40) by Sampling-based Optimization
Solving (40) is challenging, as i) the gradient of the cost function is quite difficult to obtain (Li et al., 2023), and ii) it requires global search, as evidenced by the nonconvex cost landscape in Figure 17. As a result, we resort to a simple sampling-based search which samples from the feasible set of (40) by rejection sampling, then chooses the best sample.
@@ 8.1 Roadmap Construction
1 Input: MPC Parameters: $T$, $n_{\mathrm{max}}$ and $H$;
@@ 9.1.1 Data Synthesis and Augmentation for IL
Recent advances in large-scale imitation learning are impressive (Black et al., 2024; TRI et al., 2025), but their performance hinges on large-scale, embodiment-specific datasets. This data is typically collected via human teleoperation, which presents a significant bottleneck for two reasons: not only is teleoperation time-consuming and expensive, teleoperation interfaces are also often limited to controlling robot end effectors, making it challenging for humans to provide high-quality demonstrations for tasks involving whole-body or in-hand contact.
@@ 9.1.2 Efficient Exploration for RL
Our global planner can also serve as efficient exploration mechanisms for RL. Despite many algorithmic advances, efficient exploration still remains a significant challenge for RL. For problems in Figure 16, for example, a naive Gaussian sampling exploration scheme would have a really hard time finding an effective policy. In order to bypass this difficulty, many practical RL algorithms leverage expert demonstrations either as rewards (Rajeswaran et al., 2018; Peng et al., 2018) or as random resets during training (Song et al., 2023) to guide RL exploration towards relevant states that are more beneficial to accomplish the task.
@@ 9.2 Concluding Remarks and Limitations
Nevertheless, the tools presented in this paper already enable capabilities that were previously out of reach for model-based methods. By accounting for the unilateral nature of contact, our contact trust region makes it possible to apply a broad range of robotics algorithms to contact-rich manipulation problems. We hope the MPC, grasp synthesis, and roadmap-based global planning methods introduced here are only a small sample of the many contact-rich manipulation algorithms yet to come. {acks} We thank Stephen Proulx, Xinpei Ni, and Velin Dimitrov for building the cube with active OptiTrack markers, and Giovanni Remigi for making the project website.
@@ A.4 Effect of Planning Horizon for Local MPC
The results are summarized in Table 10. Surprisingly, we found that the planning horizon $T$ has little to no effect on the final error for MPC: our method is able to reach the goals within a tight tolerance for all values of $T$. We believe that this is due to two reasons: i) we have chosen the initial configurations and goals to be reasonably reachable using our trust region, and ii) the CQDC model allows large change in configurations within a single step, so we are able to get away with significantly less horizon compared to other MPC tasks such as dynamic locomotion.
@@ A.5 Hyperparameters for Local MPC
In Table 11, we summarize additional hyperparameters for local MPC. These parameters are used for all experiments involving Algorithm 2 and Algorithm 3, including the results in Figure 11 and Figure 12.

### v4

@@ 1 Introduction
The apparent discrepancy between these two approaches raises the question of how they are related. In fact, analytically deriving the gradients from differentiable simulators shows that the Taylor expansions they produce and the complementarity constraints ubiquitous in contact-implicit trajectory optimization are fundamentally connected (Pang et al., 2023, Example 5): the Taylor expansions are obtained by applying the Implicit Function Theorem to the force-balance and complementarity conditions, which form a subset of the contact dynamics constraints typically included in CITO formulations (Fiacco, 1983; Giorgi and Zuccotti, 2018; Le Cleac’h et al., 2023).
@@
To alleviate these limitations, smoothing-based methods relax contact dynamics by reasoning about the average result of all the contact modes around a nominal state given some distribution of state and action disturbances. Recent works (Suh et al., 2022b, a) have attributed the success and scalability of Reinforcement Learning (RL) to this mechanism. Surprisingly, however, Pang et al. (2023); Zhang et al. (2023); Howell et al. (2023a) show that for contact dynamics simulated by convex optimization, this process can be done without Monte-Carlo by using barrier smoothing which performs interior-point relaxation of the original contact dynamics and produces force from a distance. However, despite occasional successes, naively applying classical control methods such as LQR on smoothed linearization of contact dynamics often leads to catastrophic failures (Shirai et al., 2025).
@@
We show the efficacy of the proposed CTR and MPC method on two representative contact-rich manipulation problems in Figure 1: (i) whole-body manipulation on the bimanual iiwa and cylinder system, and (ii) in-hand reorientation on the Allegro hand and cube system. For both systems, we aim to answer the following two questions: (i) is the MPC method successful in planning for goals under quasidynamic equations of motion (Section 5), and (ii) can the MPC method successfully stabilize quasidynamic plans under true second-order dynamics (Section 6)? Through $2000$ trajectory runs in simulation and $100$ runs on hardware, our results delve into the efficacy and the limitations of our approach, as well as the benefits of CTR over the ellipsoidal one in the context of planning and control.
@@ 2.5 Smoothing of Contact Dynamics
The resulting solution is equivalent to solving the log-barrier relaxation of (7),
@@
We note that this force (i) has a direction opposing the movement in the tangential plane, (ii) has a magnitude that satisfies (9), and iii) lies in the friction cone. In addition, plugging (12) into (11) yields
@@ 3.1 Linear Model over Smoothed Primal and Dual Variables
However, many previous works (Suh et al., 2022a; Howell et al., 2023a; Pang et al., 2023; Shirai et al., 2025; Posa et al., 2014) have noted that due to the non-smooth nature of contact, it is beneficial to smooth the dynamics where a first-order Taylor approximation would be valid beyond the contact mode to which $(\bar{q},\bar{u})$ belongs (Pang et al., 2023). A first-order Taylor expansion under such a smoothed dynamics model can be written as
@@
If we expand the Taylor-approximated equations and drop the terms above first-order, we are left with (i) the original equality conditions over nominal coordinates which must hold due to the nominal values being obtained at optimality and (ii) its first-order expansion, which must hold due to the definition of the sensitivity gradients (15). ∎
@@ 4.1 Trajectory Optimization with R-CTR
1Input: Initial state $q_{0}$, goal state $q_{\text{goal}}$, input trajectory guess $\bar{u}_{0:T-1}$, iterations limit $n_{\mathrm{max}}$;
@@ 4.3 Examples
Revisiting the 1D sphere-box system in Figure 3, we would like to push the box to a goal position $q^{\mathrm{o}}_{\text{goal}}$ by running Algorithm 1. We begin with the nominal configuration $\bar{q}_{0}=(\bar{q}_{0}^{\mathrm{o}},\bar{q}_{0}^{\mathrm{a}})=(-0.02,0.2)$, where the ball nearly touches the box, and a nominal action $^{0}\bar{u}=\bar{q}^{\mathrm{a}}_{0}$. The nominal action [^3] We use the left superscript to denote how the nominal values change through (i) applying the initial guess heuristics, and (ii) iterations in Algorithm 1. The left superscript is intended to be different from the right subscript which indexes time steps. is depicted by the yellow circle in Figure 6a. Even with the help of dynamics smoothing, the slope at this point is still small and thus provides limited guidance for planning.
@@ 4.4 Model Predictive Control
1Input: Initial state $q_{0}$, goal state $q_{\text{goal}}$, planning horizon $T$, iterations limit $n_{\mathrm{max}}$, MPC rollout horizon $H$;
@@ 6 Stabilization under Second-Order Dynamics
In this section, we would like to understand how the gap between the CQDC and real physics affect MPC performance in more realistic settings. Specifically, we focus on answering the question: can the MPC controller using the CQDC dynamics perform closed-loop stabilization (i) on high-fidelity simulated second-order dynamics, and (ii) on hardware?
@@ 6.1 Implementation Heuristics for Second-order Dynamics
Firstly, instead of running MPC as quickly as possible, we compute a longer trajectory (usually 1-2 seconds) using Algorithm 2 for multiple steps ($H>1$) (Algorithm 3), roll out $L_{u}$ in its entirety on the second-order dynamics (Algorithm 3), and repeat. This resembles the “action chuncking” scheme ubiquitous in robot learning, which has been shown to reduce compounding errors during rollouts when the policy (or controller), rather than the open-loop-stable dynamics of the robot-object system, is the primary contributor to compounding errors (Zhang et al., 2025). In our case, the MPC controller indeed injects compounding errors due to the gap between the CQDC model used for planning and the second-order nature of $f_{\text{2nd}}$. See Section A.6 for more details about the effect of $H$ on closed-loop tracking performance.
@@
1Input: Initial state $q_{0}$, goal state $q_{\text{goal}}$, planning horizon $T$, iterations limit $n_{\mathrm{max}}$, MPC rollout horizon $H$, number of re-plans NN;
@@ 7 Generating Goal-conditioned Contact Configurations
To efficiently conduct long-horizon planning for contact-rich manipulation, we utilize two key insights. The first insight comes from dynamic programming: knowledge of key intermediate states allows us to decompose a difficult long-horizon problem into multiple shorter subproblems in which greedy strategies can be more effective. For the planar-pushing example in Figure 16, consider the configuration in which the ball is at the bottom of the box as opposed to the top. If the planner had (i) seen this configuration before and (ii) known that this configuration is significantly more advantageous for pushing the box upward using local control, it can attempt to go towards this advantageous configuration first before attempting to establish contact with the box.
@@ 7.1 Problem Specification
where (40b) are joint-limit constraints, (40c) enforce non-penetration constraints for every collision pair indexed by $i$, and (40a) is our cost criteria for judging how fit $q^{\mathrm{a}}$ is in driving $q^{\mathrm{o}}$} to $q^{\mathrm{o}}_{g}$. Our cost $C$ consists of two terms: the finite-horizon value function of the MPC policy (Section 7.1.1), and a regularization term for robustness (Section 7.1.2).
@@ 7.1.2 Robustness Regularizer
To evaluate this quantity, we first make a polytopic approximation of the wrench set $\mathcal{W}_{\mathbf{\Sigma},\kappa}^{\mathcal{A}}$ by sampling from the RA-CTR using a rejection sampling scheme. This scheme (i) first samples from the ETR ellipsoid and (ii) rejects samples that do not obey feasibility constraints (this is the same procedure which generated the samples in Figure 5). Then, we fit a convex hull to these samples using the Qhull library.
@@ 7.2 Solving (40) by Sampling-based Optimization
Solving (40) is challenging, as (i) the gradient of the cost function is quite difficult to obtain (Li et al., 2023), and (ii) it requires global search, as evidenced by the nonconvex cost landscape in Figure 17. As a result, we resort to a simple sampling-based search which samples from the feasible set of (40) by rejection sampling, then chooses the best sample.
@@ 8.1 Roadmap Construction
1Input: MPC Parameters: $T$, $n_{\mathrm{max}}$ and $H$;
@@ 9.1.1 Data Synthesis and Augmentation for IL
Recent advances in large-scale imitation learning are impressive (Black et al., 2024; LBM Team TRI, 2025), but their performance hinges on large-scale, embodiment-specific datasets. This data is typically collected via human teleoperation, which presents a significant bottleneck for two reasons: not only is teleoperation time-consuming and expensive, teleoperation interfaces are also often limited to controlling robot end effectors, making it challenging for humans to provide high-quality demonstrations for tasks involving whole-body or in-hand contact.
@@ 9.1.2 Efficient Exploration for RL
Our global planner can also serve as efficient exploration mechanisms for RL. Despite many algorithmic advances, efficient exploration still remains a significant challenge for RL. For problems in Figure 16, for example, a naive Gaussian sampling exploration scheme would have a really hard time finding an effective policy. In order to bypass this difficulty, many practical RL algorithms leverage expert demonstrations either as rewards (Rajeswaran et al., 2018; Peng et al., 2018, 2021) or as random resets during training (Song et al., 2023) to guide RL exploration towards relevant states that are more beneficial to accomplish the task.
@@ 9.2 Concluding Remarks and Limitations
Nevertheless, the tools presented in this paper already enable capabilities that were previously out of reach for model-based methods. By accounting for the unilateral nature of contact, our contact trust region makes it possible to apply a broad range of robotics algorithms to contact-rich manipulation problems. We hope the MPC, grasp synthesis, and roadmap-based global planning methods introduced here are only a small sample of the many contact-rich manipulation algorithms yet to come. {acks} We thank Stephen Proulx, Xinpei Ni, and Velin Dimitrov for building the cube with active OptiTrack markers, Giovanni Remigi for making the project website, and Bridget Hogan for building visualization tools.
{dci}
The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article
{funding}
@@ A.4 Effect of Planning Horizon for Local MPC
The results are summarized in Table 10. Surprisingly, we found that the planning horizon $T$ has little to no effect on the final error for MPC: our method is able to reach the goals within a tight tolerance for all values of $T$, even $T=1$.
We believe this can be explained by the Bellman’s principle of optimality. A 1-step greedy policy is long-term optimal if it minimizes the sum of the immediate 1-step cost and the true optimal cost-to-go. This has been experimentally verified on smooth dynamical systems (Orrico et al., 2024).
When casting our MPC cost (36a) for $T=1$ into the Bellman equation, $\|u_{1}-u_{0}\|^{2}$ corresponds to the 1-step cost, and $\|q_{\text{goal}}-q_{1}\|^{2}_{\mathbf{Q}}$ serves as an approximation of the cost-to-go. The empirical success of $T=1$ suggests that for goals generated by our method (Section A.3), the simple quadratic term is a highly effective approximation of the true optimal cost-to-go.
@@ A.5 Hyperparameters for Local MPC
In Table 11, we summarize additional hyperparameters for local MPC. These parameters are kept the same for all experiments involving Algorithm 2 and Algorithm 3, including the results in Figure 11 and Figure 12.